OpenStudy (anonymous):

Given: The equation of a line y= -5x+2. Part a. Write the equation of the line in slope-intercept form parallel to the given line that contains the point (3, -7) Part b. Write the equation of a line in slope-intercept form perpendicular to the given line that contains the point (10, -7).

OpenStudy (agent47):

Slope of the original line: -5. a)The lines are parallel, so the slopes are the same. Form the equation: y-y1 = -5(x-x1)

OpenStudy (agent47):

y - (-7) = -5(x-3). Simplify.

OpenStudy (anonymous):

y=-5x+22?

OpenStudy (texaschic101):

slope is -5. Parallel line will have same slope. y = mx + b slope(m) = -5 (3,-7) x = 3 and y = -7 now we sub and solve for b -7 = -5(3) + b -7 = -15 + b -7 + 15 = b 8 = b your parallel equation is : y = -5x + 8 ======================== slope is -5. Perpendicular line will have negative reciprocal slope. All that means is " flip " the slope and change the sign. Negative reciprocal of -5, or -5/1, is 1/5. y = mx + b slope(m) = 1/5 (10,-7) x = 10 and y = -7 now we sub -7 = 1/5(10) + b -7 = 2 + b -7 - 2 = b -9 = b perpendicular equation is : y = 1/5x - 9 ======================== any questions ?

OpenStudy (anonymous):

Nope, thank you :)